# Robust Statistics vs. Machine Learning vs. Bayesian Inference: Insights into Handling Faulty GNSS Measurements in Field Robotics

### v1

@@ Abstract
This paper presents research findings on handling faulty measurements (i.e., outliers) of global navigation satellite systems (GNSS) for robot localization under adverse signal conditions in field applications, where raw GNSS data are frequently corrupted due to environmental interference such as multipath, signal blockage, or non-line-of-sight conditions. In this context, we investigate three strategies applied specifically to GNSS pseudorange observations: robust statistics for error mitigation, machine learning for faulty measurement prediction, and Bayesian inference for noise distribution approximation.
@@
Our goal is to examine both established techniques and newly proposed methods, thereby advancing the understanding of how to handle faulty range measurements, such as GNSS, for robust, long-term robot localization. In addition to presenting successful results, this work highlights critical observations and open questions to motivate future research in robust state estimation.
@@ I Motivation
Robot navigation in large-scale field environments commonly depends on Global Navigation Satellite Systems (GNSS) to maintain globally consistent state estimates. However, in complex environments such as urban canyons and forest, GNSS signals are frequently degraded by multipath propagation and non-line-of-sight (NLOS) conditions. These effects lead to measurement outliers with time-varying noise characteristics that are often heavy-tailed, skewed, and multimodal, making them difficult to model a priori [1].
Consequently, state estimators based on least-squares techniques may degrade or even diverge when relying on inconsistent measurement models with overconfident, hand-crafted noise parameters, as they oftentimes underestimate the complexity and magnitude of uncertainty. As a result, developing a robot localization approach that ensures robust long-term operation across diverse environments remains a significant challenge.
@@ I-A3 Robust Error Modeling
Limitations: Most studies using non-adaptive M-estimators focus on hard-redescending kernels like Huber and Tukey, overlooking the potential of soft-redescending functions for robust GNSS-based robot localization. They also lack analysis of parameter tuning and the trade-off between robustness and efficiency. Although some works [24, 25, 26] evaluate various kernels and settings, they are not tailored to range-based localization.
@@
These gaps highlight the need for a comprehensive study of M-estimators across categories, focusing on their mathematical properties and suitability for robust GNSS-based robot localization.
@@ I-A5 Learning-based Methods
As environment-dependent outliers may be challenging to identify and can exhibit complex noise patterns that are difficult to model with traditional statistical methods, learning-based methods111Although many learning-based methods are rooted in the statistical inference and thus share similar theoretical foundations, we intentionally categorize inference-based and learning-based methods as distinct approaches. have also gained attention for handling faulty GNSS observations.
@@
Moreover, as there is currently no evidence that learning-based methods cannot generalize across varying scenarios in vehicle localization, further investigation is needed into how learning concepts can be effectively integrated to enhance the robustness of state estimation applications.
@@ I-B Objective
II Preminilaries
II-A GNSS Pseudorange Model for Robot Localization
@@
where the vectors and represent the positions of the GNSS antenna and -th satellite in the Earth-centered, Earth-fixed (ECEF) frame, respectively. The variables and represent the bias due to receiver clock delay and satellite clock delay. The tropospheric, ionospheric and multipath delays are denoted as , and . The pseudorange noise is .
@@ III-B Error Mitigation using M-Estimators
The kernel function in (3) is typically designed to reduce the influence of a large residual , which are often caused by outliers. Numerous kernel functions222Due to space limitations, we do not provide detailed technical explanations and instead refer readers to [24, 25]. are available in M-estimation, typically classified by their influence functions into three categories: robust-monotonic, robust-soft-redescending, and robust-hard-redescending.
[…]
@@ III-C Machine Learning for Outlier Prediction
To address the limitations of current learning-based approaches for faulty GNSS prediction (see Section I-A5), we take into account the noise characteristics of GNSS observations in urban environments and proposed a transformer-enhanced Long short-term Memory (TE-LSTM) network in our previous work [35].
We study the proposed network alongside a baseline LSTM network [34] and two classical machine learning models: support vector machine (SVM) and extreme gradient boosting (XGBoost), aiming to provide an comprehensive analysis comparing classical models with deep networks, as well as evaluating performance on in-distribution and out-of-distribution data.
@@ III-D Bayesian Inference for Noise Distribution Approximation
[…]
@@
In this context, we observed that two key issues arise from this. The first issue arises whenever the robot is in motion. As it moves relative to a mostly static environment, its position with respect to reflective surfaces changes, leading to varying residuals caused by the same structures. Although the Gaussian mixture model includes a component with a reasonable mean, it does not account for this relative positional change. Second, if the environment changes abruptly, the outdated GMM trained with past residuals may fail to capture new outliers, leading to poor state estimates. This creates a feedback loop: degraded estimates lead to inaccurate residuals, which further distort the GMM and worsen the next estimate.
@@
with , ensuring 99% of lies within probablistic bounds.
@@
with denoting the covariance of the Cauchy hypothesis, and , representing the mean and precision of the dominant GMM mode. The discrete selector determines the active noise model and is jointly optimized with continuous variables based on the following pseudo-probablity
@@ IV-A Data Specification
We conduct data from real-world measurement compagins to train and test the invistigated methods introduced above. Table II summarized the essential data characteristics.
@@ IV-B4 Discussions
However, we do not propose a universal strategy for tuning hyper-parameters, as generalizing M-estimation for long-term, large-scale localization remains challenging. Outlier characteristics and environmental variability create a difficult trade-off between robustness and efficiency in GNSS-based localization, where measurement corruption often coincides with signal loss. Even well-tuned estimators may fail to ensure reliable performance across diverse scenarios.
@@
[…]
@@ IV-C4 Discussion
[…]
@@ IV-D Handling Outliers using Variational Bayesian Inference
Discussions: However, although the proposed inference method, updating only the noise models, can significantly improve the overall performance of robot localization, issues such as the smoothness of the estimated trajectory remain unresolved. Fig.,7 exemplifies this issue, showing that the estimator using MH-GMM exhibits high variations in 2-D errors due to unsmooth estimates. This may be caused by inaccurate estimation of the switching variable , which results in frequent transitions between different noise models.
@@ V Conclusion
Building on the results of this research, our objective is to explore various techniques for handling measurement outliers in range observations to enable robust robot localization in the field. Our methods, experimental evaluations, limitations, and open questions are summarized in this paper, providing a foundation for future work in enabling resilient robot localization in challenging field scenarios.

### v2

# Robust Statistics vs. Machine Learning vs. Bayesian Inference: Insights into Handling Faulty GNSS Measurements for Robust Vehicle Localization

@@ Abstract
This paper presents research findings on handling faulty measurements (i.e., outliers) of global navigation satellite systems (GNSS) for vehicle localization under adverse signal conditions in field applications, where raw GNSS data are frequently corrupted due to environmental interference such as multipath, signal blockage, or non-line-of-sight conditions. In this context, we investigate three strategies applied specifically to GNSS pseudorange observations: robust statistics for error mitigation, machine learning for faulty measurement prediction, and Bayesian inference for noise distribution approximation.
@@
Our goal is to examine both established techniques and newly proposed methods, thereby advancing the understanding of how to handle faulty range measurements, such as GNSS, for robust, long-term vehicle localization. In addition to presenting successful results, this work highlights critical observations and open questions to motivate future research in robust state estimation.
@@ I Motivation
Vehicle navigation in large-scale field environments commonly depends on Global Navigation Satellite Systems (GNSS) to maintain globally consistent state estimates. However, in complex environments such as urban canyons and forest, GNSS signals are frequently degraded by multipath propagation and non-line-of-sight (NLOS) conditions. These effects lead to measurement outliers with time-varying noise characteristics that are often heavy-tailed, skewed, and multimodal, making them difficult to model a priori [1].
Consequently, state estimators based on least-squares techniques may degrade or even diverge when relying on inconsistent measurement models with overconfident, hand-crafted noise parameters, as they oftentimes underestimate the complexity and magnitude of uncertainty. As a result, developing a vehicle localization approach that ensures robust long-term operation across diverse environments remains a significant challenge.
@@ I-A3 Robust Error Modeling
Limitations: Most studies using non-adaptive M-estimators focus on hard-redescending kernels like Huber and Tukey, overlooking the potential of soft-redescending functions for robust GNSS-based vehicle localization. They also lack analysis of parameter tuning and the trade-off between robustness and efficiency. Although some works [24, 25, 26] evaluate various kernels and settings, they are not tailored to range-based localization.
@@
These gaps highlight the need for a comprehensive study of M-estimators across categories, focusing on their mathematical properties and suitability for robust GNSS-based vehicle localization.
@@ I-A5 Learning-based Methods
As environment-dependent outliers may be challenging to identify and can exhibit complex noise patterns that are difficult to model with traditional statistical methods, learning-based methods111Although many learning-based methods are rooted in statistical inference and thus share similar theoretical foundations, we intentionally categorize inference-based and learning-based methods as distinct approaches. have also gained attention for handling faulty GNSS observations.
@@
Moreover, as there is currently no evidence that learning-based methods can generalize across varying scenarios in vehicle localization, further investigation is needed into how learning concepts can be effectively integrated to enhance the robustness of state estimation applications.
@@ I-B Objective
II Preliminaries
II-A GNSS Pseudorange Model for Vehicle Localization
@@
where the vectors and represent the positions of the GNSS antenna and -th satellite in the Earth-centered, Earth-fixed (ECEF) frame, respectively. The variables and represent the bias due to receiver clock delay and satellite clock delay. The tropospheric, ionospheric and multipath delays are denoted as , , and . The pseudorange noise is .
@@ III-B Error Mitigation using M-Estimators
The kernel function in (3) is typically designed to reduce the influence of a large residual , which is often caused by outliers. Numerous kernel functions222Due to space limitations, we do not provide detailed technical explanations and instead refer readers to [24, 25]. are available in M-estimation, typically classified by their influence functions into three categories: robust-monotonic, robust-soft-redescending, and robust-hard-redescending.
Two key concepts, robustness and relative efficiency, are used to evaluate M-estimators, with a trade-off noted by [37]: “The more robust an estimator is, the less efficient it is”. Consequently, relative efficiency serves as a fundamental reference for tuning M-estimators.
@@ III-C Machine Learning for Outlier Prediction
To address the limitations of current learning-based approaches for faulty GNSS prediction (see Section I-A5), we take into account the noise characteristics of GNSS observations in urban environments and propose a transformer-enhanced Long Short-Term Memory (TE-LSTM) network in our previous work [35].
We study the proposed network alongside a baseline LSTM network [34] and two classical machine learning models: support vector machine (SVM) and extreme gradient boosting (XGBoost), aiming to provide a comprehensive analysis comparing classical models with deep networks, as well as evaluating performance on in-distribution and out-of-distribution data.
@@ III-D Bayesian Inference for Noise Distribution Approximation
Fig. 6 illustrates the concept of the proposed algorithm within a nested estimation framework. We employ multi-sensor factor graph optimization (FGO) to obtain “near-optimal” state estimates, which are then used to compute the pseudorange residuals. In a parallel stage, these residuals are sampled over a fixed time window, and a variational expectation-maximization (EM) algorithm is applied to approximate a GMM for each satellite [38, Chapter 10].
@@
In this context, we observed that two key issues arise from this. The first issue arises whenever the vehicle is in motion. As it moves relative to a mostly static environment, its position with respect to reflective surfaces changes, leading to varying residuals caused by the same structures. Although the Gaussian mixture model includes a component with a reasonable mean, it does not account for this relative positional change. Second, if the environment changes abruptly, the outdated GMM trained with past residuals may fail to capture new outliers, leading to poor state estimates. This creates a feedback loop: degraded estimates lead to inaccurate residuals, which further distort the GMM and worsen the next estimate.
@@
with , ensuring 99% of lies within probabilistic bounds.
@@
with denoting the covariance of the Cauchy hypothesis, and , representing the mean and precision of the dominant GMM mode. The discrete selector determines the active noise model and is jointly optimized with continuous variables based on the following pseudo-probability
@@ IV-A Data Specification
We conduct data from real-world measurement campaigns to train and test the investigated methods introduced above. Table II summarizes the essential data characteristics.
@@ IV-B4 Discussions
However, we do not propose a universal strategy for tuning hyperparameters, as generalizing M-estimation for long-term, large-scale localization remains challenging. Outlier characteristics and environmental variability create a difficult trade-off between robustness and efficiency in GNSS-based localization, where measurement corruption often coincides with signal loss. Even well-tuned estimators may fail to ensure reliable performance across diverse scenarios.
@@
This motivates a new open question: “Can we predict measurement outliers using learning-based methods before integrating them into the estimator?” Or more generally: “Can we leverage learning to recover the information needed for tuning M-estimators?”
@@ IV-C4 Discussion
In addition, discriminative modeling with pre-trained models relies on a global model assumed to generalize across applications. This raises a key question: “Does a truly global model for outlier characteristics exist?”
Put differently: “Do we need a global model at all?” Or can robustness be enhanced through a “local” model that is updated in time and online? This leads us to the next method.
@@ IV-D Handling Outliers using Variational Bayesian Inference
Discussions: However, although the proposed inference method, updating only the noise models, can significantly improve the overall performance of vehicle localization, issues such as the smoothness of the estimated trajectory remain unresolved. Fig. 7 exemplifies this issue, showing that the estimator using MH-GMM exhibits high variations in 2-D errors due to unsmooth estimates. This may be caused by an inaccurate estimation of the switching variable , which results in frequent transitions between different noise models.
@@ V Conclusion
Building on the results of this research, our objective is to explore various techniques for handling measurement outliers in range observations to enable robust vehicle localization in the field. Our methods, experimental evaluations, limitations, and open questions are summarized in this paper, providing a foundation for future work in enabling resilient vehicle localization in challenging field scenarios.